Let z be a normal random variable with a mean of 35 and a standard deviation of 4. Determine (z ≤ 1.4)

1
Expert's answer
2022-03-29T11:02:35-0400

"P(z<1.4)=P(z<\frac{x-\mu}{\sigma})=P(z<\frac{1.4-35}{4})=P(z<-8.425)=0"